# Information Inequalities for Five Random Variables

### v1

@@ 5.1. Case I
$\begin{array}[]{llllll}\mathsf{C}_{1}:&[abcd\mkern 1.5mu],\\[3.0pt] \mathsf{C}_{2}\mbox{--}\mathsf{C}_{4}:&(a,b\mkern 1.0mu{|}\mkern 1.0muc),&(a,c\mkern 1.0mu{|}\mkern 1.0mub),&(b,c\mkern 1.0mu{|}\mkern 1.0mua),\\[3.0pt] \mathsf{C}_{5}\mbox{--}\mathsf{C}_{7}:&(a,b\mkern 1.0mu{|}\mkern 1.0mud),&(a,d\mkern 1.0mu{|}\mkern 1.0mub),&(b,d\mkern 1.0mu{|}\mkern 1.0mua),\\[3.0pt] \mathsf{C}_{8}\mbox{--}\mathsf{C}_{11}:&(c,d\mkern 1.0mu{|}\mkern 1.0mua),&(c,d|b),&(c,d),&(a,b\mkern 1.0mu{|}\mkern 1.0mucd),\\[3.0pt] \mathsf{C}_{12}\mbox{--}\mathsf{C}_{14}:&(a,b\mkern 1.0mu{|}\mkern 1.0muz),&(a,z\mkern 1.0mu{|}\mkern 1.0mub),&(b,z\mkern 1.0mu{|}\mkern 1.0mua),\\[3.0pt] \mathsf{C}_{15}\mbox{--}\mathsf{C}_{19}:&(a\mkern 1.0mu{|}\mkern 1.0mubcd),&(b\mkern 1.0mu{|}\mkern 1.0muacd),&(c\mkern 1.0mu{|}\mkern 1.0muabd),&(d\mkern 1.0mu{|}\mkern 1.0muabc),&(z\mkern 1.0mu{|}\mkern 1.0muab).\end{array}$
Coordinates $\mathsf{C}_{15}\mbox{--}\mathsf{C}_{19}$ cover the modular part of $\mathcal{Q}^{*}$. The tight part is spanned by the coordinate vectors $\mathsf{C}_{1}\mbox{--}\mathsf{C}_{14}$, each of these vectors is orthogonal to the modular part. Let $\tilde{P}_{1}$ be the inverse of the matrix of this coordinate transformation, and the vector $\mathbf{p}_{1}$ be the first row of $\tilde{P}_{1}$. Let $P_{1}$ be the submatrix formed from rows 221414 of $\tilde{P}$. Coordinates of the vector $\mathbf{x}\in\mathbb{R}^{19}$ in this coordinate system are $\tilde{P}_{1}\mkern 1.5mu\mathbf{x}^{\top}$, and, in particular, the Ingleton value $f[abcd\mkern 1.5mu]$ is the scalar product $\mathbf{p}_{1}\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{x}$. Consequently, the tight part of the intersection of $\mathcal{Q}^{*}$ and the hyperplane $abcd\mkern 1.5mu]=-1$ in this coordinate system is
@@
$n$ rows columns    facets vertices time (sec)
@@
Table 1 shows the sizes of the generating matrix $M$, the total number of facets and vertices (including extremal directions) of the cross-section $\mathcal{Q}^{*}_{1}$, and the running time of the vertex enumeration algorithm on a single-core desktop computer with an Intel® Core i5-4590 CPU @ 3.30GHz processor and 8GB of memory. The running time was taken up almost exclusively by the underlying LP solver. While the number of facets grows quite moderately with $n$, the number of vertices more than doubles at each generation. The matrix $M$, despite the numerous improvements, is highly degenerate, and numerical instability, originating from both the LP solver and the applied polyhedral algorithm, prevented the completion of the computation for larger values of $n$. The results of the computation are presented in Section 6.
@@ 5.2. Case II
The cone $\mathcal{Q}^{*}$ is intersected with the hyperplane $acbd\mkern 1.5mu]=-1$. The coordinate system is similar to the one used in Case I. Base elements $b$ and $c$ are swapped in coordinates $\mathsf{C}_{1}\mbox{--}\mathsf{C}_{11}$, while the other coordinates remain unchanged. The tight part of the intersection, denoted by $\mathcal{Q}^{*}_{2}$, is defined similarly with the same matrix $M$ but a different coordinate transformation matrix $\tilde{P}_{2}$, vector $\mathbf{p}_{2}$, and submatrix $P_{2}$ as
@@
$n$ rows columns    facets vertices time (sec)
@@
The running time is significantly shorter than in Case I. It is explained by the fact that the polyhedral algorithm requires solving an LP instance for each vertex and each facet in the result, and those numbers are significantly smaller here. The generating matrix $M$ is the same in both cases, implying that the problem size is the same. Numerical instability prevented completing the computation for $n=10$ even in this case.
@@ 5.3. Case III
Changing the first $11$ coordinates of the coordinate system used in Case I to
$\begin{array}[]{llllll}\mathsf{C}_{1}:&[cdab\mkern 1.5mu],\\[3.0pt] \mathsf{C}_{2}\mbox{--}\mathsf{C}_{4}:&(c,d\mkern 1.0mu{|}\mkern 1.0mua,&(a,c\mkern 1.0mu{|}\mkern 1.0mud),&(a,d\mkern 1.0mu{|}\mkern 1.0muc),\\[3.0pt] \mathsf{C}_{5}\mbox{--}\mathsf{C}_{7}:&(b,c\mkern 1.0mu{|}\mkern 1.0mud),&(c,d\mkern 1.0mu{|}\mkern 1.0mub),&(b,d\mkern 1.0mu{|}\mkern 1.0muc),\\[3.0pt] \mathsf{C}_{8}\mbox{--}\mathsf{C}_{11}:&(a,b\mkern 1.0mu{|}\mkern 1.0muc),&(a,b|d),&(a,b),&(c,d\mkern 1.0mu{|}\mkern 1.0muab),\end{array}$
and keeping the rest, the vertex enumeration algorithm used in the previous cases generated the vertices and extremal directions of the $13$-dimensional tight part of $\mathcal{Q}^{*}_{3}$. The computation showed that it is a pointed cone with a single vertex that has coordinates $\mathsf{C}_{2}\mbox{--}\mathsf{C}_{14}$ equal to zero (while $\mathsf{C}_{1}=-1$); and has $14$ extremal directions, $12$ of which are coordinate axes. Polymatroids representing the extremal directions are linear when restricted to the base set abcdabcd (they satisfy $f[cdab\mkern 1.5mu]=0$, therefore, the other Ingleton values are also non-negative). Consequently, these polymatroids have an $n$-copy for all $n\geqslant 1$. Finally, the remaining polymatroid at the single vertex has $f(a,b)=0$ (as the coordinate $\mathsf{C}_{10}$ is zero), which means that $f\mathord{\upharpoonright}ab$ is modular. By Claim 6 it also has an $n$-copy for all $n\geqslant 1$. This concludes the proof that no non-Shannon inequality is generated in this case.
@@ 6. Experimental Information Inequalities
For a fixed $n\geqslant 1$, the problem of extracting the set of non-Shannon inequalities that form the necessary and sufficient conditions for the existence of an $n$-copy of a polymatroid on the base set abcdzabcdz was shown to be equivalent to determining all facets of a $13$-dimensional polyhedral cone. The cone was cut into several pieces and the facets of each piece were computed for $n\leqslant 9$. In this section we take a quick look at the computational results. Below the symbols $\mathcal{Z}$, $\mathcal{C}$, $\mathcal{D}$ denote the following entropy expressions:
$\mathcal{Z}$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(a,z\mkern 1.0mu{|}\mkern 1.0mub)+(b,z\mkern 1.0mu{|}\mkern 1.0mua),$
$\mathcal{C}$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(a,c\mkern 1.0mu{|}\mkern 1.0mub)+(b,c\mkern 1.0mu{|}\mkern 1.0mua),$
$\mathcal{D}$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(a,d\mkern 1.0mu{|}\mkern 1.0mub)+(b,d\mkern 1.0mu{|}\mkern 1.0mua).$
@@ 6.1. Case I
In the $abcd\mkern 1.5mu]<0$ case, facets of the polyhedron $\mathcal{Q}^{*}_{1}$ from (5) include all $13$ coordinate planes orthogonal to the coordinate axes $\mathsf{C}_{2}$$\mathsf{C}_{14}$. These facets correspond to the non-negativity of the expression defining the coordinate. $\mathcal{Q}^{*}_{1}$ has two additional Shannon facets, corresponding to the Shannon inequalities $(a,z)\geqslant 0$ and $(b,z)\geqslant 0$. The remaining facets determine the non-Shannon inequalities we are interested in. They come in three flavors:
$(a,b\mkern 1.0mu{|}\mkern 1.0muz)+\alpha_{s}[abcd\mkern 1.5mu]+\alpha_{s}\mathcal{Z}+\beta_{s}\mathcal{C}+\gamma_{s}\mathcal{D}$ $\geqslant 0,$ (7)
$(a,b\mkern 1.0mu{|}\mkern 1.0muc)+\alpha_{s}[abcd\mkern 1.5mu]+\,(\alpha_{s}\mkern 1.5mu{+}\mkern 1.5mu\beta_{s})\mathcal{C}\mkern 1.5mu\,+\gamma_{s}\mathcal{D}$ $\geqslant 0,$ (8)
$(a,b\mkern 1.0mu{|}\mkern 1.0mud)+\alpha_{s}[abcd\mkern 1.5mu]+\,\beta_{s}\mathcal{C}\mkern 1.5mu\,+(\alpha_{s}\mkern 1.5mu{+}\mkern 1.5mu\gamma_{s})\mathcal{D}$ $\geqslant 0,$ (9)
where $\langle\alpha_{s},\beta_{s},\gamma_{s}\rangle$ are certain triplets of non-negative integers. For illustration, consider the $n=3$ case. As reported in Table 1, for $n=3$ the polyhedron $Q^{*}_{1}$ has $34$ facets. These facets determine $15$ Shannon inequalities, $11$ inequalities of the form (7), and $44$ inequalities of the form (8) and (9). The $\langle\alpha,\beta,\gamma\rangle$ triplets appearing in (7) are listed in three columns in Table 3. Inequalities in (8) and (9) use triplets from the first column only; these are the triplets that also appear in the $n=2$ generation.
$\alpha$ $\beta$ $\gamma$
1 0 0
2 0 1
2 1 0
3 1 1
$\alpha$ $\beta$ $\gamma$
3 0 3
3 3 0
4 1 3
4 3 1
$\alpha$ $\beta$ $\gamma$
6 3 5
6 5 3
7 5 5
Table 3. Coefficient values for $n=3$
In general, inequalities in (8) and (9) are consequences of (7) via replacing $z$ with $c$ and $d$, respectively. Since the copy $c_{n}$ of $c$ in an $n$-copy polymatroid $f^{*}$ can be considered to be the variable $z$ in the $n\mkern 1.5mu{-}\mkern 1.5mu1$-copy when $f^{*}$ is restricted to $N^{*}\mkern 1.5mu{-}\mkern 1.5mu\{d_{n}z\}$, inequalities valid for $n\mkern 1.5mu{-}\mkern 1.5mu1$-copy instances must hold in an $n$-copy with $z$ replaced by $c$, and, similarly, when $z$ is replaced by $d$. This property is confirmed by the computational results. Additionally, all inequalities not containing the variable $z$ proved to be derivatives from the previous generation via the above substitutions. The main goal in Section 7 and Section 8 is to obtain a general description of the triplets $\langle\alpha_{s},\beta_{s},\gamma_{s}\rangle$ occurring in (7).
@@ 6.2. Case II
Inequalities in the $acbd\mkern 1.5mu]<0$ case have a similar but significantly simpler structure. Facets of the $n$-copy cone $\mathcal{Q}^{*}_{2}$ from (6) include the coordinate planes, the two Shannon facets $(a,z)\geqslant 0$ and $(b,z)\geqslant 0$ as above, and additional facets generating the inequalities
$(a,b\mkern 1.0mu{|}\mkern 1.0muz)+k\mkern 1.5mu[acbd\mkern 1.5mu]+k\mkern 1.5mu\mathcal{Z}+\frac{k(k\mkern 1.5mu{-}\mkern 1.5mu1)}{2}\mathcal{C}$ $\geqslant 0,$
$(a,b\mkern 1.0mu{|}\mkern 1.0muc)+k\mkern 1.5mu[acbd\mkern 1.5mu]+\frac{(k\mkern 1.5mu{+}\mkern 1.5mu1)k}{2}\mathcal{C}$ $\geqslant 0$
for $1\leqslant k\leqslant n$ for the first, and $1\leqslant k\leqslant n\mkern 1.5mu{-}\mkern 1.5mu1$ for the second set of inequalities. As noted in the $abcd\mkern 1.5mu]<0$ case, inequalities in the second set are instances of ones from the first from the previous generation when $z$ is replaced by $c$. When $z$ is replaced by $d$, the resulting inequality
$(a,b\mkern 1.0mu{|}\mkern 1.0mud)+k\mkern 1.5mu[acbd\mkern 1.5mu]+k\mkern 1.5mu\mathcal{D}+\frac{k(k\mkern 1.5mu{-}\mkern 1.5mu1)}{2}\mathcal{C}\geqslant 0$
is Shannon as $acbd\mkern 1.5mu]+\mathcal{D}\geqslant 0$ holds in every polymatroid.

### v2

@@ 5.1. Case I
$\begin{array}[]{llllll}\mathsf{C}_{1}:&[abcd\mkern 1.5mu],\\[3.0pt] \mathsf{C}_{2}\mbox{--}\mathsf{C}_{4}:&(a,b\mkern 1.5mu{|}\mkern 1.5muc),&(a,c\mkern 1.5mu{|}\mkern 1.5mub),&(b,c\mkern 1.5mu{|}\mkern 1.5mua),\\[3.0pt] \mathsf{C}_{5}\mbox{--}\mathsf{C}_{7}:&(a,b\mkern 1.5mu{|}\mkern 1.5mud),&(a,d\mkern 1.5mu{|}\mkern 1.5mub),&(b,d\mkern 1.5mu{|}\mkern 1.5mua),\\[3.0pt] \mathsf{C}_{8}\mbox{--}\mathsf{C}_{11}:&(c,d\mkern 1.5mu{|}\mkern 1.5mua),&(c,d|b),&(c,d),&(a,b\mkern 1.5mu{|}\mkern 1.5mucd),\\[3.0pt] \mathsf{C}_{12}\mbox{--}\mathsf{C}_{14}:&(a,b\mkern 1.5mu{|}\mkern 1.5muz),&(a,z\mkern 1.5mu{|}\mkern 1.5mub),&(b,z\mkern 1.5mu{|}\mkern 1.5mua),\\[3.0pt] \mathsf{C}_{15}\mbox{--}\mathsf{C}_{19}:&(a\mkern 1.5mu{|}\mkern 1.5mubcd),&(b\mkern 1.5mu{|}\mkern 1.5muacd),&(c\mkern 1.5mu{|}\mkern 1.5muabd),&(d\mkern 1.5mu{|}\mkern 1.5muabc),&(z\mkern 1.5mu{|}\mkern 1.5muab).\end{array}$
Coordinates $\mathsf{C}_{15}\mbox{--}\mathsf{C}_{19}$ cover the modular part of $\mathcal{Q}^{*}$. The tight part is spanned by the coordinate vectors $\mathsf{C}_{1}\mbox{--}\mathsf{C}_{14}$, and each of these vectors is orthogonal to the modular part. Let $\tilde{P}_{1}$ be the inverse of the matrix of this coordinate transformation, and the vector $\mathbf{p}_{1}$ be the first row of $\tilde{P}_{1}$. Let $P_{1}$ be the submatrix formed from rows $2$ to $14$ of $\tilde{P}$. Coordinates of the vector $\mathbf{x}\in\mathbb{R}^{19}$ in this coordinate system are $\tilde{P}_{1}\mkern 1.5mu\mathbf{x}^{\top}$, and, in particular, the Ingleton value $f[abcd\mkern 1.5mu]$ is the scalar product $\mathbf{p}_{1}\mkern 1.5mu{\cdot}\mkern 1.5mu\mathbf{x}$. Consequently, the tight part of the intersection of $\mathcal{Q}^{*}$ and the hyperplane $abcd\mkern 1.5mu]=-1$ in this coordinate system is
@@
$n$ Rows Columns    Facets Vertices Time (s)
@@
Table 1 shows the sizes of the generating matrix $M$, the total number of facets and vertices (including extremal directions) of the cross-section $\mathcal{Q}^{*}_{1}$, and the running time of the vertex enumeration algorithm on a single-core desktop computer with an Intel® Core™ i5-4590 CPU @ 3.30GHz processor and 8 GB of memory. The running time was taken up almost exclusively by the underlying LP solver. The number of columns is the dimension $d_{3}$ from formula (5). While the number of facets grows quite moderately with $n$, the number of vertices more than doubles at each generation. The matrix $M$, despite numerous improvements, is highly degenerate, and numerical instability, originating from both the LP solver and the applied polyhedral algorithm, prevented the completion of the computation for larger values of $n$. The results of the computation are presented in Section 6.
@@ 5.2. Case II
The cone $\mathcal{Q}^{*}$ is intersected with the hyperplane $acbd\mkern 1.5mu]=-1$. The coordinate system is similar to the one used in Section 5.1. Base elements $b$ and $c$ are swapped in coordinates $\mathsf{C}_{1}$ to $\mathsf{C}_{11}$, while the other coordinates remain unchanged. The tight part of the intersection, denoted by $\mathcal{Q}^{*}_{2}$, is defined similarly with the same matrix $M$ but a different coordinate transformation matrix $\tilde{P}_{2}$, vector $\mathbf{p}_{2}$, and submatrix $P_{2}$ as
@@
$n$ Rows Columns    Facets Vertices Time (s)
@@
The running time is significantly shorter than in Section 5.1. It is explained by the fact that the polyhedral algorithm requires solving an LP instance for each vertex and each facet in the result, and those numbers are significantly smaller here. The generating matrix $M$ is the same in both cases, implying that the problem size is the same. Numerical instability prevented completing the computation for $n=10$ even in this case.
@@ 5.3. Case III
Changing the first $11$ coordinates of the coordinate system used in Section 5.1 to
$\begin{array}[]{llllll}\mathsf{C}_{1}:&[cdab\mkern 1.5mu],\\[3.0pt] \mathsf{C}_{2}\mbox{--}\mathsf{C}_{4}:&(c,d\mkern 1.5mu{|}\mkern 1.5mua,&(a,c\mkern 1.5mu{|}\mkern 1.5mud),&(a,d\mkern 1.5mu{|}\mkern 1.5muc),\\[3.0pt] \mathsf{C}_{5}\mbox{--}\mathsf{C}_{7}:&(b,c\mkern 1.5mu{|}\mkern 1.5mud),&(c,d\mkern 1.5mu{|}\mkern 1.5mub),&(b,d\mkern 1.5mu{|}\mkern 1.5muc),\\[3.0pt] \mathsf{C}_{8}\mbox{--}\mathsf{C}_{11}:&(a,b\mkern 1.5mu{|}\mkern 1.5muc),&(a,b|d),&(a,b),&(c,d\mkern 1.5mu{|}\mkern 1.5muab),\end{array}$
and keeping the rest, the vertex enumeration algorithm used in the previous cases generated the vertices and extremal directions of the $13$-dimensional tight part of $\mathcal{Q}^{*}_{3}$. The computation showed that it is a pointed cone with a single vertex that has coordinates $\mathsf{C}_{2}\mbox{--}\mathsf{C}_{14}$ equal to zero (while $\mathsf{C}_{1}=-1$) and has $14$ extremal directions, $12$ of which are coordinate axes. Polymatroids representing the extremal directions are linear when restricted to the base set abcdabcd (they satisfy $f[cdab\mkern 1.5mu]=0$; therefore, the other Ingleton values are also non-negative). Consequently, these polymatroids have an $n$-copy for all $n\geqslant 1$. Finally, the remaining polymatroid at the single vertex has $f(a,b)=0$ (as the coordinate $\mathsf{C}_{10}$ is zero), which means that $f\mathord{\upharpoonright}ab$ is modular. By Claim 6 it also has an $n$-copy for all $n\geqslant 1$. This concludes the proof that no non-Shannon inequality is generated in this case.
@@ 6. Experimental Information Inequalities
For a fixed $n\geqslant 1$, the problem of extracting the set of non-Shannon inequalities that form the necessary and sufficient conditions for the existence of an $n$-copy of a polymatroid on the base set abcdzabcdz was shown to be equivalent to determining all facets of a $14$-dimensional polyhedral cone. The cone was cut into several pieces and the facets of each piece were computed for $n\leqslant 9$. In this section we take a quick look at the computational results. In the description the symbols $\mathcal{Z}$, $\mathcal{C}$, $\mathcal{D}$ denote the following entropy expressions:
$\mathcal{Z}$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(a,z\mkern 1.5mu{|}\mkern 1.5mub)+(b,z\mkern 1.5mu{|}\mkern 1.5mua),$
$\mathcal{C}$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(a,c\mkern 1.5mu{|}\mkern 1.5mub)+(b,c\mkern 1.5mu{|}\mkern 1.5mua),$
$\mathcal{D}$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(a,d\mkern 1.5mu{|}\mkern 1.5mub)+(b,d\mkern 1.5mu{|}\mkern 1.5mua).$
@@ 6.1. Case I
In the $abcd\mkern 1.5mu]<0$ case, facets of the polyhedron $\mathcal{Q}^{*}_{1}$ from (6) include all $13$ coordinate planes orthogonal to the coordinate axes $\mathsf{C}_{2}$$\mathsf{C}_{14}$. These facets correspond to the non-negativity of the expression defining that coordinate. $\mathcal{Q}^{*}_{1}$ has two additional Shannon facets, corresponding to the Shannon inequalities $(a,z)\geqslant 0$ and $(b,z)\geqslant 0$. The remaining facets determine the non-Shannon inequalities we are interested in. They come in three flavors:
$(a,b\mkern 1.5mu{|}\mkern 1.5muz)+\alpha_{s}[abcd\mkern 1.5mu]+\alpha_{s}\mathcal{Z}+\beta_{s}\mathcal{C}+\gamma_{s}\mathcal{D}$ $\geqslant 0,$ (8)
$(a,b\mkern 1.5mu{|}\mkern 1.5muc)+\alpha_{s}[abcd\mkern 1.5mu]+\,(\alpha_{s}\mkern 1.5mu{+}\mkern 1.5mu\beta_{s})\mathcal{C}\mkern 1.5mu\,+\gamma_{s}\mathcal{D}$ $\geqslant 0,$ (9)
$(a,b\mkern 1.5mu{|}\mkern 1.5mud)+\alpha_{s}[abcd\mkern 1.5mu]+\,\beta_{s}\mathcal{C}\mkern 1.5mu\,+(\alpha_{s}\mkern 1.5mu{+}\mkern 1.5mu\gamma_{s})\mathcal{D}$ $\geqslant 0,$ (10)
where $\langle\alpha_{s},\beta_{s},\gamma_{s}\rangle$ are certain triplets of non-negative integers. For illustration, we consider the $n=3$ case. As reported in Table 1, for $n=3$ the polyhedron $Q^{*}_{1}$ has $34$ facets. These facets determine $15$ Shannon inequalities (among which there are $13$ coordinate planes), $11$ inequalities of the form (8), and $44$ inequalities of the form (9) and (10). The $\langle\alpha,\beta,\gamma\rangle$ triplets appearing in (8) are listed in three columns in Table 3. Inequalities in (9) and (10) use triplets from the first column only; these are the triplets that also appear in the $n=2$ generation.
$\alpha,\beta,\gamma$ $\alpha,\beta,\gamma$ $\alpha,\beta,\gamma$
$\langle 1,0,0\rangle$ $\langle 3,0,3\rangle$ $\langle 6,3,5\rangle$
$\langle 2,0,1\rangle$ $\langle 3,3,0\rangle$ $\langle 6,5,3\rangle$
$\langle 2,1,0\rangle$ $\langle 4,1,3\rangle$ $\langle 7,5,5\rangle$
$\langle 3,1,1\rangle$ $\langle 4,3.1\rangle$
Table 3. The $\alpha,\beta,\gamma$ coefficient values for $n=3$
In general, inequalities in (9) and (10) are consequences of (8) by replacing $z$ with $c$ and $d$, respectively. Since the copy $c_{n}$ of $c$ in an $n$-copy polymatroid $f^{*}$ can be considered to be the variable $z$ in the $n\mkern 1.5mu{-}\mkern 1.5mu1$-copy when $f^{*}$ is restricted to $N^{*}\mkern 1.5mu{-}\mkern 1.5mu\{d_{n}z\}$, inequalities valid for $n\mkern 1.5mu{-}\mkern 1.5mu1$-copy instances must hold in an $n$-copy with $z$ replaced by $c$, and, similarly, when $z$ is replaced by $d$. This property is confirmed by the computational results. Additionally, all inequalities not containing the variable $z$ proved to be derivatives from the previous generation via the above substitutions. The main goal of Sections 7 and 8 is to obtain a general description of the triplets $\langle\alpha_{s},\beta_{s},\gamma_{s}\rangle$ occurring in (8).
@@ 6.2. Case II
Inequalities in the $acbd\mkern 1.5mu]<0$ case have a similar but significantly simpler structure. Facets of the $n$-copy cone $\mathcal{Q}^{*}_{2}$ from (7) include $13$ coordinate planes, the two Shannon facets $(a,z)\geqslant 0$ and $(b,z)\geqslant 0$ as above, and additional facets corresponding to the inequalities
$(a,b\mkern 1.5mu{|}\mkern 1.5muz)+k\mkern 1.5mu[acbd\mkern 1.5mu]+k\mkern 1.5mu\mathcal{Z}+\frac{k(k\mkern 1.5mu{-}\mkern 1.5mu1)}{2}\mathcal{C}$ $\geqslant 0~~~~~\mbox{ where }1\leqslant k\leqslant n,$
$(a,b\mkern 1.5mu{|}\mkern 1.5muc)+k\mkern 1.5mu[acbd\mkern 1.5mu]+\frac{(k\mkern 1.5mu{+}\mkern 1.5mu1)k}{2}\mathcal{C}$ $\geqslant 0~~~~~\mbox{ where }1\leqslant k<n.$
As noted in the $abcd\mkern 1.5mu]<0$ case, inequalities in the second set are instances of ones from the first from the previous generation when $z$ is replaced by $c$. When $z$ is replaced by $d$, the resulting inequality
$(a,b\mkern 1.5mu{|}\mkern 1.5mud)+k\mkern 1.5mu[acbd\mkern 1.5mu]+k\mkern 1.5mu\mathcal{D}+\frac{k(k\mkern 1.5mu{-}\mkern 1.5mu1)}{2}\mathcal{C}\geqslant 0$
is Shannon as $acbd\mkern 1.5mu]+\mathcal{D}=(a,d\mkern 1.5mu{|}\mkern 1.5muc)+2(b,d\mkern 1.5mu{|}\mkern 1.5mua)+(a,c\mkern 1.5mu{|}\mkern 1.5mub)\geqslant 0$ holds in every polymatroid.